Entries for Treasury Stock

On May 27, Kick Off Inc. reacquired 6,300 shares of its common stock at $50 per share. On August 3, Kick Off sold 3,000 of the reacquired shares at $53 per share. November 14, Kick Off sold the remaining shares at $49 per share.

Journalize the transactions of May 27, August 3, and November 14. For a compound transaction, if an amount box does not require an entry, leave it blank.

May 27
Aug. 3
Nov. 14
Answer #1
May 27 Treasury stock 315000 =6300*50
      Cash 315000
Aug 3 Cash 159000 =3000*53
      Treasury stock 150000 =3000*50
      Paid in Capital from sale of Treasury stock 9000
Nov 14 Cash 161700 =3300*49
Paid in Capital from sale of Treasury stock 3300
      Treasury stock 165000 =3300*50